# How to Make the Most of LLMs’ Grammatical Knowledge for Acceptability Judgments

### v1

@@ 4 Experimental Setup
We use six LLMs, among which Llama-3-70B, Mixtral-8x7B-v0.1, and Qwen2-57B-A14B are base models, while Llama-3-70B-Instruct, Mixtral-8x7B-Instruct-v0.1, and Qwen2-57B-A14B-Instruct are instruct models based on the pre-trained counterparts. We hereafter abbreviate these models, e.g., to Llama-3, omitting the model sizes and minor versions. Post-training for the three instruct models includes supervised fine-tuning on an instruction dataset, i.e., instruction-tuning, and aforementioned DPO Meta (2024); Jiang et al. (2024); Team (2024). They are models publicly available on Hugging Face Hub. All six models are used for English experiments, while Chinese experiments examine only Qwen2 and Qwen2-Instruct, which are trained on Chinese texts. On inference, we perform 4-bit quantization using bitsandbytes777 https://github.com/TimDettmers/bitsandbytes to compress the models.888 See Section B.1 for the computational budgets.
@@
We use two MP acceptability judgment benchmarks: BLiMP Warstadt et al. (2020) for English and CLiMP Xiang et al. (2021) for Chinese. BLiMP is composed of minimal pairs from 67 different paradigms, each containing 1,000 pairs of sentences. The paradigms are grouped into 12 categories of linguistic phenomena. CLiMP consists of 16 paradigms, each with 1,000 pairs like BLiMP. Its paradigms are categorized into 9 linguistic phenomena. The linguistic phenomena and licenses of the benchmarks are detailed in Section A.1 and Section A.2, respectively.
@@
We evaluate the methods by accuracy. Random chance accuracy is 50%, as the task is a binary classification one.
5 Results
Table 3 summarizes the results. The statistics of the in-template probability readout methods and prompting methods are the average of the five scores by the five versions of templates. They show that Yes/No prob comp and in-template LP are particularly strong in both languages. On BLiMP, Yes/No prob comp achieves the highest mean accuracy for five out of six models. The mean accuracies of Llama-3-Instruct and Qwen2 exceed that of humans (the majority vote of 20 crowd workers) reported in Warstadt et al. (2020), 88.6%. In-template LP scores in the top two methods for accuracy on all but one model. The results on CLiMP find similar conclusions; the top two methods are in-template LP and Yes/No prompting.
Sentence readout methods, LP, MeanLP, and PenLP, which have been dominant in previous studies, underperformed in-template LP for all settings. This indicates that including some guidance about the task in the input to LLMs improves acceptability judgment performance.
Methods giving two sentences to the model, i.e., A/B prompting and in-template comparative LP, also underperformed Yes/No prob comp and in-template LP, respectively, though giving multiple choices is a common approach to harness LLMs’ knowledge for a classification task as seen in studies such as Hendrycks et al. (2021).999 See Section C.2 for an analysis of the underperformance of A/B prompting.
6 Analysis
Given the excellence of Yes/No prob comp and in-template LP, this section provides further analysis to reveal their strengths and weaknesses.
@@ 6 Analysis
Figure 2: Major methods’ correlation between the token-length difference ($|\bm{s}_{\mathrm{acceptable}}|-|\bm{s}_{\mathrm{unacceptable}}|$) and the accuracy (best template) by model. The shadow denotes 95% confidence intervals.
@@
These observations are quantitatively supported by the correlation coefficient between the token-length difference and the dichotomous variable that gets 1 for a successful prediction and 0 for a failure. Table 4 shows the average coefficients101010 The point biserial correlation coefficient is mathematically equivalent to the Pearson correlation coefficient. of Yes/No prob comp are much closer to zero than those of readout methods on both benchmarks, demonstrating its robustness against the token-length bias. This, in turn, indicates that the readout methods need better normalization techniques.
@@
In-template LP -0.118 -0.147
In-template PenLP -0.094 -0.269
Yes/No prob comp -0.019 -0.006
Table 4: Major methods’ point biserial correlation coefficient between the prediction success and token-length difference (averaged over models) by benchmark. The bold font denotes the value closest to zero.
(a) BLiMP
(b) CLiMP
Figure 3: Accuracy of major methods (best template) and humans, by linguistic phenomenon and model. For BLiMP, phenomena where either method wins the other by at least 1% for at least five models are underlined.
@@
Llama-3 Llama-3 -Instruct Mixtral Mixtral -Instruct Qwen2 Qwen2 -Instruct Qwen2 Qwen2 -Instruct
Ensemble P-only 76.0 89.0 85.4 84.4 89.5 87.0 86.6 85.0
Ensemble Mix-P3 82.7 $\mathbf{89.7}$ $\mathbf{87.5}$ $\mathbf{86.6}$ $\mathbf{90.2}$ $\mathbf{87.7}$ 87.5 87.5
Ensemble Mix-L3 $\mathbf{86.1}$ 86.3 86.5 86.4 86.7 84.8 $\mathbf{89.7}$ $\mathbf{88.6}$
Ensemble L-only 85.1 84.1 84.6 84.3 84.3 81.0 88.4 87.0
In-template LP (oracle) 85.0 84.2 84.6 84.5 84.1 81.3 88.2 86.5
Yes/No prob comp (oracle) 77.8 89.3 85.6 87.5 89.2 87.4 86.6 85.2
@@
Figure 3 illustrates the accuracy of in-template LP, Yes/No prob comp, and the humans by linguistic phenomenon; the scores of humans are from Warstadt et al. (2020) and Xiang et al. (2021). For in-template LP and Yes/No prob comp, the result of the best-performing template is shown. Here we find that the two methods have different strengths. On BLiMP, Yes/No prob comp excels at phenomena such as Subject-verb agreement (S-v agr.) and Binding for most (at least five out of six) models.111111 See Section A.1 for examples of these phenomena. In contrast, in-template LP is superior in Ellipsis and Quantifiers for most models. On CLiMP, Yes/No prob comp is good at Coverb and in-template LP at NP head finality (NP head). This indicates that each method harnesses different aspects of the models’ grammatical knowledge.
Given the aforementioned token-length bias, one hypothesis to explain this difference would be that Yes/No prob comp is stronger in phenomena with large token-length differences. Our analysis on BLiMP, however, does not support this. See Section C.3 for the details.
Meanwhile, some phenomena are challenging for both methods. As Figure 3 shows, on BLiMP, the two methods underperform humans for all models in Island effects and Quantifiers, which were shown to be challenging also by Warstadt et al. (2020). On CLiMP, our methods struggle with phenomena such as Binding and Passive, lagging far behind human performance.
@@
Now we have 10 sets of predictions by the two methods, as each has five templates. To compare ensembling single-method predictions and ensembling multi-method predictions on equal terms, we sample five without replacement from the 10 and perform majority voting by the five. We prepare the following four settings, which differ in the balance between the two methods: P-only, Mix-P3, Mix-L3, and L-only. P-only and L-only are ensembles of predictions by Yes/No prob comp only and in-template LP only, respectively. Mix-P3 and Mix-L3 use three predictions from Yes/No prob comp and in-template LP, respectively, with two predictions from the other method. We report the mean accuracy of 10 trials for these settings because the result is non-deterministic due to sampling.
Table 5 demonstrates that ensembles of the two methods, either Mix-P3 or Mix-L3, yield the best results across models, surpassing the oracle (max) accuracies of methods without ensembling, except for Mixtral-Instruct. The highest score by Mix-P3 with Qwen2 is 1.6 points higher than humans (described in Section 5). This indicates that the two methods have complementary capabilities.
Figure 4: Major methods’ accuracy (best template) on S-v agr. paradigms in BLiMP by attractor type and model.
Figure 5: Accuracy on BLiMP by paradigm type, method (row), and model (column). Each dot represents a paradigm. Short horizontal bars denote the mean accuracy of the category. The dashed lines denote the chance accuracy. Results by the best template are used for each method.
@@
Attractors refer to material intervening agreement dependencies, and their effects on acceptability judgments have been studied. Below are examples of different attractor types in S-v agr., from Warstadt et al. (2020); (a) contains no attractor, (b) has an attractor as a relational noun, and (c) has an attractor in a relative clause.121212 Subject-verb agreement does not exist in Chinese, so we focus on English here.
@@
Word-shuffling paradigms are challenging.
BLiMP’s 67 paradigms can be divided into two categories based on whether minimal-pair sentences of the paradigm have the same bag of words when the cases are ignored. We call the paradigms where this is true word-shuffling paradigms.131313 CLiMP does not have word-shuffling paradigms. Following is an example pair from a word-shuffling paradigm, existential_there_quantifiers_2.
@@
    Each book is there disturbing Margaret.141414 This sentence is non-sensical, which could lower the accuracy of judgments.
@@
    *There is each book disturbing Margaret.
Figure 5 shows the accuracy by paradigm, paradigm type—word-shuffling or not, method, and model, demonstrating that the word-shuffling paradigms have much lower accuracy than other phenomena across methods and models. The accuracy of word-shuffling paradigms averaged over models and methods is 71.6% compared to 87.9% of other paradigms. The paradigm marking the lowest accuracy is aforementioned existential_there_quantifiers_2, whose accuracy is only 39.9% on average. Note that such a large difference is not observed for humans according to the data by Warstadt et al. (2020); humans’ accuracy on word-shuffling paradigms and other paradigms are, on average, 83.1% and 89.7%, respectively. This suggests that word-shuffling paradigms remain a challenge for the current LLMs, as they have trouble recognizing word shuffling that corrupts grammar even with our best-performing methods.
7 Conclusion
To investigate how best to measure LLMs’ grammatical knowledge, we compared nine acceptability judgment methods across six LLMs and two languages. We found that in-template LP and Yes/No prob comp consistently outperform conventional sentence probability readout methods. This indicates that sentence probability readout methods are suboptimal and should be replaced in an evaluation of LLMs. Meanwhile, the performant two methods excel in different phenomena, suggesting they harness different aspects of LLMs’ grammatical knowledge. We thus recommend using diverse judgment methods for a more comprehensive and appropriate evaluation of LLMs.
8 Limitations
One of this paper’s key findings is that in-template LP and Yes/No prob comp excel in different linguistic phenomena. To investigate the reasons for the differences, we examined hypotheses that Yes/No prob comp is stronger in phenomena where the acceptable sentence is, on average, longer than the unacceptable one (See Section C.3). Yet the hypotheses were not supported, leaving the cause of their different strengths an open question.
@@ 8 Limitations
Appendix A Benchmarks
A.1 Linguistic Phenomena
@@ A.2 URLs and Licenses
Appendix B Experiments
B.1 Computational Budgets
For each method or combination of methods and templates, we used a single NVIDIA A6000 GPU with 48GB RAM. The total GPU hours are estimated to be about 126 hours and 7 hours for the BLiMP and CLiMP experiments, respectively.
Appendix C Results and Analysis
C.1 Max Accuracy

### v2

@@ 4 Experimental Setup
We use eight state-of-the-art LLMs, among which Llama-3-70B Meta (2024), Mixtral-8x7B-v0.1 Jiang et al. (2024), Qwen2-57B-A14B Qwen Team (2024), and Yi-1.5-34B Young et al. (2024) are base models, while Llama-3-70B-Instruct, Mixtral-8x7B-Instruct-v0.1, Qwen2-57B-A14B-Instruct, and Yi-1.5-34B-Chat are instruct models based on the pre-trained counterparts. These models ranked relatively high in the leaderboard of English language understanding888 https://paperswithcode.com/sota/multi-task-language-understanding-on-mmlu or Chinese LLMs999 https://huggingface.co/spaces/BAAI/open_cn_llm_leaderboard at the time of model selection. For Chinese experiments, we substituted Yi-1.5 for Mixtral because Mixtral is not explicitly trained for Chinese tasks, while Yi-1.5 ranked high in the Chinese LLM leaderboard. We hereafter abbreviate these models, e.g., to Llama-3, omitting the model sizes and minor versions. Post-training for the three instruct models includes supervised fine-tuning on an instruction dataset, i.e., instruction-tuning and DPO. They are publicly available on Hugging Face Hub. On inference, we perform 4-bit quantization using bitsandbytes101010 https://github.com/TimDettmers/bitsandbytes to compress the models. The computational budgets are described in Appendix C.
@@
We use two MP acceptability judgment benchmarks: BLiMP Warstadt et al. (2020) for English and CLiMP Xiang et al. (2021) for Chinese. BLiMP is composed of minimal pairs from 67 different paradigms, each containing 1,000 pairs of sentences. The paradigms are grouped into 12 categories of linguistic phenomena. CLiMP consists of 16 paradigms, each with 1,000 pairs like BLiMP. The paradigms are grouped into 9 linguistic phenomena. We focus on these two because no other MP benchmarks contain hundreds or thousands of sentences for each paradigm, to our knowledge, which is important for reliable experiments. The linguistic phenomena and licenses of the two benchmarks are detailed in Section B.1 and Section B.2, respectively.
@@
We evaluate the methods by accuracy. Random chance accuracy is 50%, as the two classes are balanced in our benchmarks.
@@
Figure 2: Top methods’ correlation between the token-length difference ($|\bm{s}_{\mathrm{acceptable}}|-|\bm{s}_{\mathrm{unacceptable}}|$) and the accuracy (best template) by model, showing the robustness of Yes/No prob comp. The shadow denotes 95% confidence intervals.
5 Results
Table 3 summarizes the results. The statistics of the in-template probability readout methods and prompting methods are the average of the five scores by the five versions of templates.111111 We conducted paired bootstrap resampling Koehn (2004) to validate whether each of our methods statistically significantly improves accuracy compared to LP, the best-performing conventional method; we performed 1,000 resamplings with replacement, sampling 67,000 and 16,000 instances each time for BliMP and CLiMP, respectively.
Comparison between methods reveals the effectiveness of in-template LP and Yes/No prob comp. In-template LP achieves significantly higher accuracies than LP in all settings, i.e., benchmark-model pairs, across languages. On CLiMP, it marks the highest accuracy for all models. Sentence probability readout methods—LP, MeanLP, and PenLP—underperformed in-template LP in all settings, although they have been dominant in previous studies. This indicates that including guidance about the task in the input to LLMs improves judgment performance. Meanwhile, Yes/No prob comp achieves the highest accuracy for five out of six models on BLiMP; the mean accuracies of Llama-3-Instruct and Qwen2 exceed that of humans (the majority vote of 20 crowd workers) reported in Warstadt et al. (2020), 88.6%.
Methods giving two sentences to the model—A/B prompting and in-template comparative LP—consistently underperformed Yes/No prob comp and in-template LP, respectively, suggesting that LLMs struggle to handle choice identifiers such as A and B (See Section D.2 for more analysis, which reveals the low performance of A/B prompting can be partly attributed to a preference for a specific choice). This may suggest that making LLMs select an identifier from multiple choices—a common approach for classification tasks such as question answering Hendrycks et al. (2021)—is suboptimal for making full use of LLMs’ knowledge in general; substituting our choice-free methods could improve performance in other tasks.
6 Analysis
Given the remarkable performance of Yes/No prob comp and in-template LP, this section further investigates their strengths and weaknesses.
(a) BLiMP
(b) CLiMP
Figure 3: Accuracy of top methods (best template) and humans, by linguistic phenomenon and model, demonstrating the difference in strengths between methods. For each benchmark, phenomena where either method wins the other by at least 1 point for at least five models are underlined.
@@ 6 Analysis
These observations are quantitatively supported by the correlation coefficient between the token-length difference and the dichotomous variable that gets 1 for a successful prediction and 0 for a failure. Table 4 shows the average coefficients of Yes/No prob comp are much closer to zero than those of readout methods on both benchmarks, demonstrating its robustness against the token-length bias. This, in turn, indicates that the readout methods need better normalization techniques.
@@
In-template LP $-0.118$ $-0.135$
In-template PenLP --0.094 --0.182
Yes/No prob comp $\mathbf{-0.019}$-$\mathbf{0.051}$
Table 4: Top methods’ point biserial correlation coefficient between the prediction success and token-length difference (averaged over models) by benchmark. The bold font denotes the value closest to zero.
@@
Llama-3 Llama-3 -Instruct Qwen2 Qwen2 -Instruct Llama-3 Llama-3 -Instruct Qwen2 Qwen2 -Instruct
Ensemble L0:P5 76.0 89.0 89.5 87.0 76.0 76.6 87.6 84.1
Ensemble L2:P3 82.7 $\mathbf{89.7}$ $\mathbf{90.2}$ $\mathbf{87.7}$ 79.1 78.4 89.4 87.1
Ensemble L3:P2 $\mathbf{86.1}$ 86.3 86.7 84.8 $\mathbf{87.1}$ $\mathbf{83.8}$ $\mathbf{90.5}$ $\mathbf{89.4}$
Ensemble L5:P0 85.1 84.1 84.3 81.0 86.1 83.0 88.4 87.0
In-template LP (oracle) 85.0 84.2 84.1 81.3 86.1 83.2 88.2 86.5
Yes/No prob comp (oracle) 77.8 89.3 89.2 87.4 78.1 78.4 87.5 84.3
Table 5: Percentage accuracy of voting ensembles of in-template LP and Yes/No prob comp, with the oracle (max) accuracy by single methods (best template). The bold font denotes the best ensemble score. Results on Mixtral and Yi models are omitted because the same trend is observed; see Section D.4 for their results.
Figure 4: Accuracy on BLiMP by paradigm type, method, and model (best template), showing the difficulty of word-shuffling paradigms. Each dot represents a paradigm, short bars denote the mean accuracy of the category, cross markers denote mean human accuracies, and dashed lines denote the chance accuracy.
@@
Figure 3 illustrates the accuracy of in-template LP, Yes/No prob comp, and the humans by linguistic phenomenon; the scores of humans are from Warstadt et al. (2020) and Xiang et al. (2021). For in-template LP and Yes/No prob comp, the result of the best-performing template is shown. Here we find that the two methods have different strengths. On BLiMP, Yes/No prob comp excels at phenomena such as Subject-verb agreement (S-v agr.) and Binding for most (at least five out of six) models (See Section B.1 for examples of these phenomena). In contrast, in-template LP is superior in Ellipsis and Quantifiers for most models. On CLiMP, Yes/No prob comp is good at Coverb and in-template LP at NP head finality (NP head). This indicates that each method harnesses different aspects of the models’ grammatical knowledge.
Given the aforementioned token-length bias, one hypothesis to explain this difference would be that Yes/No prob comp is more accurate in phenomena with large token-length differences. Our analysis on BLiMP, however, does not support this, as detailed in Section D.3.
Meanwhile, some phenomena are challenging for both methods. As Figure 3 shows, on BLiMP, the two methods underperform humans for all models in Island effects and Quantifiers, which were shown to be challenging also by Warstadt et al. (2020). On CLiMP, our methods struggle with phenomena such as Binding and Passive.
@@
Now we have 10 sets of predictions by the two methods, as each has five templates. To compare ensembling single-method predictions and ensembling multi-method predictions on equal terms, we sample five without replacement from the 10 and perform majority voting by the five. We prepare the following four settings, which differ in the balance between the two methods: P-only, Mix-P3, Mix-L3, and L-only. P-only and L-only are ensembles of predictions by Yes/No prob comp only and in-template LP only, respectively. Mix-P3 and Mix-L3 use three predictions from Yes/No prob comp and in-template LP, respectively, with two predictions from the other method. We report the mean accuracy of 10 trials for these settings as the result is non-deterministic due to sampling.
Table 5 demonstrates that ensembles of the two methods, either Mix-P3 or Mix-L3, yield the best results across models, surpassing the oracle (max) accuracies of methods without ensembling, except for Mixtral-Instruct on BLiMP. The highest score by Mix-P3 with Qwen2 is 1.6 points higher than humans (described in Section 5). This indicates that the two methods have complementary capabilities.
Figure 5: Accuracy on S-v agr. paradigms in BLiMP by attractor type, method, and model (best template), showing that attractors in a relative clause hinder acceptability judgments.
Word-shuffling paradigms are challenging.
BLiMP’s 67 paradigms can be divided into two categories based on whether paired sentences of the paradigm have the same bag of words when the cases are ignored. We call the paradigms where this is true word-shuffling paradigms. In other words, the unacceptable sentence in word-shuffling paradigms can be obtained by shuffling the words in the acceptable counterpart. Here, we focus on BLiMP because CLiMP does not have word-shuffling paradigms. Following is an example pair from a word-shuffling paradigm, coordinate_structure_constraint_complex_left_branch.
  1. (a)
    What reports did Rose hate and James find?
  2. (b)
    *What did Rose hate reports and James find?
Figure 4 shows the accuracy by paradigm, paradigm type—word-shuffling or not, demonstrating that the word-shuffling paradigms have much lower accuracy than other phenomena across methods and models. The accuracy of word-shuffling paradigms averaged over models and methods is 71.6% compared to 87.9% of other paradigms. The best accuracy on word-shuffling paradigms, achieved by Yes/No prob comp with Qwen2, was 79.6%, which is much lower than that of the same method on other paradigms, 91.3%. Note that such large differences are not observed for humans according to the data released by Warstadt et al. (2020); humans’ accuracy on word-shuffling paradigms and other paradigms are, on average, 83.1% and 89.7%, respectively. This suggests that word-shuffling paradigms remain a challenge for the current LLMs, as they have trouble recognizing word shuffling that corrupts grammar even with our best-performing method.
Why are word-shuffling paradigms difficult? We hypothesize that the LLMs are insensitive to the word order of the inputs. Previously, Sinha et al. (2021b) demonstrated the word-order insensitivity of Transfer-based models, such as RoBERTa Liu et al. (2019) and BART Lewis et al. (2020), in the task of natural language inference. Sinha et al. (2021a) and Pham et al. (2021) report the results to the same effect. The Transformer architecture, which forms the basis of the models they and we used, may generally struggle to capture the difference in word order.
@@
Attractors refer to material intervening agreement dependencies, and previous work has shown that they can impair acceptability judgments. Below are examples of different attractor types in S-v agr., from Warstadt et al. (2020); (a) contains no attractor, (b) has an attractor as a relational noun, and (c) has an attractor in a relative clause. Because subject-verb agreement does not exist in Chinese, we focus on English here.
@@
7 Conclusion
We investigate how to derive the most accurate acceptability judgments from LLMs by comparing nine methods, using eight LLMs and two benchmarks. Our experiments reveal that in-template LP consistently outperforms conventional sentence probability readout methods while Yes/No prob comp achieves the highest accuracies on the English benchmark. Our analysis demonstrates that the two methods excel in different phenomena, suggesting they harness different aspects of LLMs’ grammatical knowledge. We also find that ensembling the two methods achieves even higher accuracy. Consequently, we recommend ensembling the two methods or employing in-template LP as more effective alternatives to conventional approaches. Meanwhile, we show that word-shuffling paradigms remain difficult for all our methods, posing a challenge for future work.
8 Limitations
In BLiMP and CLiMP, acceptable sentences are designed to be grammatical or well-formed, and the sentences in each paradigm were validated by humans Warstadt et al. (2020); Xiang et al. (2021). However, acceptable sentences can be non-sensical as they are automatically generated without their senses being considered. For example, the acceptable sentence in the following pair from the existential_there_quantifiers_2 phenomenon is non-sensical; it is difficult to identify what situation is described in the acceptable sentence, at least without context.
@@ 8 Limitations
    All convertibles weren’t there existing.
@@
    *There weren’t all convertibles existing.
To make correct acceptability judgments, our prompts or in-template inputs guided the LLMs to focus on acceptability or grammatical correctness. However, this may be insufficient when making judgments on phenomena that contain many non-sensical acceptable sentences like the above. To address this issue, future work could use prompts or in-template inputs that clearly instruct the model to focus only on the form of each sentence and ignore whether the sentence is sensical.
One of the key findings of this paper is that in-template LP and Yes/No prob comp excel in different linguistic phenomena. To investigate the reasons for the differences, we examined a hypothesis that Yes/No prob comp is more accurate in phenomena where the acceptable sentence is, on average, longer than the unacceptable one (See Section D.3). Yet the hypotheses were not supported, leaving the cause of their different strengths an open question.
@@
Additionally, this paper focused on English and Chinese acceptability judgments. Although this was inevitable to secure a sufficient data size per benchmark, experiments in more languages are desired for more robust conclusions.
Acknowledgments
We are grateful to Hongyu Sun for verifying the Chinese translations of the prompts and templates.
Appendix A Templates and Prompts
Input Type Example Input
Sentence 王玉珍震惊了她自己。
[…]
Table 6: Example Chinese inputs of the readout methods. The target or inserted sentences are underlined. See Table 1 for English versions.
Type Role Example Message
A/B System 您的任务是评估给定文本的质量。
[…]
Yes/No System 您的任务是评估给定文本的质量。
[…]
Table 7: Example Chinese messages for prompting. The target or inserted sentences are underlined. See Table 2 for English versions.
Appendix B Benchmarks
B.1 Linguistic Phenomena
@@ B.2 URLs and Licenses
Appendix C Computational Budgets
For each experiment on a method or a combination of a method and template, we used a single NVIDIA A6000 GPU with 48GB RAM. The total GPU hours are estimated to be about 126 hours and 21 hours for the BLiMP and CLiMP experiments, respectively.
Appendix D Results and Analysis
D.1 Max Accuracy
Llama-3 Llama-3-Instruct Mixtral Mixtral-Instruct Qwen2 Qwen2-Instruct
LP 79.6 77.1 82.5 82.3 80.4 79.7
MeanLP 77.1 74.8 79.6 79.4 77.7 77.1
PenLP 79.2 76.8 82.2 82.0 79.9 79.2
In-template LP $\mathbf{85.0}$ 84.2 84.6 84.5 84.1 81.3
In-template MeanLP 83.1 82.6 83.2 83.1 82.8 80.5
In-template PenLP 84.4 83.5 84.5 84.3 83.6 81.2
In-template compar. LP 76.4 66.0 75.2 69.8 67.7 63.6
A/B prompting 79.5 83.9 81.9 83.6 81.5 82.8
Yes/No prob comp 77.8 $\mathbf{89.3}$ $\mathbf{85.6}$ $\mathbf{87.5}$ $\mathbf{89.2}$ $\mathbf{87.4}$
(a) BLiMP
Llama-3 Llama-3-Instruct Yi-1.5 Yi-1.5-Chat Qwen2 Qwen2-Instruct
LP 83.2 80.4 86.8 85.3 85.4 85.4
MeanLP 74.5 71.7 75.9 74.5 74.5 74.3
PenLP 80.3 77.7 84.3 82.3 82.2 82.0
In-template LP $\mathbf{86.1}$ $\mathbf{83.2}$ $\mathbf{87.9}$ $\mathbf{87.8}$ $\mathbf{88.2}$ $\mathbf{86.5}$
In-template MeanLP 80.7 78.5 80.4 81.0 78.9 79.0
In-template PenLP 83.6 81.1 85.1 84.7 83.8 83.5
In-template compar. LP 72.2 61.6 68.2 67.3 74.2 63.5
A/B prompting 71.7 74.2 78.6 79.7 83.2 82.5
Yes/No prob comp 78.1 78.4 79.7 82.1 87.5 84.3
(a) CLiMP
Table 11: Percentage max accuracy by method and model. The bold font denotes the best scores. Underlines denote the second best.
D.2 Why A/B prompting does not perform well